# POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization

### v2

@@ Appendix A Annotation Guidelines
    Gender Polarization – Discrimination based on gender
@@
    Sexual Orientation Polarization – Tension or marginalization based on sexual orientation
  6. 2.6
@@
Note: Multiple types may be selected if applicable.
@@
    Deindividuation – Ignoring individual identities Example: “All migrants are thieves.”
@@
    Extreme Language or Absolutism – Using words like “always,” “never,” or “worst” Example: “We can never trust them.”
@@
Thank you for your valuable contribution to this research.
Appendix B Prompt for Text Classification
Task 1: Polarization Detection – Zero Shot
Given the following texts in [X] language, classify them as polarized or non-polarized based on the following definition:
Definition of Polarization: Polarization refers to the phenomenon where opinions, beliefs, or behaviors become more extreme or divided, resulting in increased separation and conflict between different groups. These groups may be based on politics or ideology, ethnicity or race, religion, gender, or sexual orientation. Polarization involves negative attitudes such as stereotyping, denigration, dehumanization, de-individualization, or intolerance toward individuals or groups outside one’s own, while exhibiting blind support and solidarity toward those within the group.
Reasoning: Provide the reason why the text demonstrates polarization.
Label: Return True if the text is polarized, otherwise False. Do not return any other value.
Task 1: Polarization Detection – Few Shot
Same instructions as above. Examples provided:
  • Text: Black immigrants should leave the USA.
    Reasoning: This statement stereotypes a group based on race and immigration status, implying a negative attitude and exclusion.
    Label: True
  • Text: Africans are lazy because they are backward.
    Reasoning: Negative trait assignment based on ethnicity, showing polarization.
    Label: True
  • Text: Muslims are terrorists.
    Reasoning: Promotes harmful stereotypes and dehumanization.
    Label: True
  • Text: African women are beautiful.
    Reasoning: Positive generalization, not polarizing.
    Label: False
Task 2: Polarized Text Typing – Zero Shot
Classify each polarized text into one or more of the following types:
  • Gender/Sexual: Discrimination based on gender or sexual orientation
  • Political: Division or conflict between political ideologies
  • Religious: Conflict between religious groups
  • Racial/Ethnic: Conflict or division based on race or ethnicity
  • Other: Related to economy, technology, media, etc.
Output: A list of five binary values in the order: [Gender/Sexual, Political, Religious, Racial/Ethnic, Other]
Few Shot Examples
  • Text: Black immigrants should leave the USA.
    Output: [0, 0, 0, 1, 0]
  • Text: Muslims are terrorists.
    Output: [0, 0, 1, 0, 0]
  • Text: Women should be confined to traditional roles.
    Output: [1, 0, 0, 0, 0]
  • Text: The economy is failing because of government policies.
    Output: [0, 1, 0, 0, 0]
Task 3: Polarization Manifestations – Zero Shot
Analyze and identify the following manifestations in each text. Output a list of six binary values in this order:
[Vilification, Extreme Language, Stereotype, Invalidation, Lack of Empathy, Dehumanization]
Definitions
  • Stereotype/Deindividuation: Generalizes traits, ignores individuality.
  • Vilification: Defames a group or person, inciting fear.
  • Dehumanization: Strips humanity, compares people to animals or objects.
  • Extreme Language: Uses absolutes like always, never, us vs. them.
  • Lack of Empathy: Dismisses others’ experiences or views.
  • Invalidation: Denies the identity or existence of a group.
Few Shot Examples
  • Text: Black immigrants should leave the USA.
    Output: [1, 1, 1, 1, 1, 0]
  • Text: Africans are lazy because they are backward.
    Output: [1, 1, 1, 1, 1, 0]
  • Text: Muslims are terrorists.
    Output: [1, 1, 1, 1, 1, 1]
  • Text: African women are beautiful.
    Output: [0, 0, 0, 0, 0, 0]
Appendix C Dataset Samples
@@ A Federal Investigation needs to be launched against TWITTER Fraud and Election Interference P EL, I
Text Lang Label
@@
Text Lang Label

### v3

@@ Appendix E Annotation Guidelines
    Gender/Sexual Polarization – Discrimination based on gender or sexual orientation
@@
Note: Select all relevant options if the text contains more than one of the above.
@@
    Extreme Language or Absolutism – Using words like “always”, “never”, or “worst”. Example: “We can never trust them.”
@@
Note: Select all relevant options if the text contains more than one of the above.
Appendix F Prompts for Text Classification
F.1 Prompts for polarization detection
TASK: Determine whether the following {{lang_full}} text is polarized or not polarized. DEFINITION: Polarization refers to the process or phenomenon in which opinions, beliefs, or behaviors become more extreme or divided, leading to a greater distance or conflict between differing groups. Attitude polarization is the negative attitude that individuals or groups display towards individuals and groups outside their group while also showing blind support and solidarity towards people within their group. Polarization denotes stereotyping, vilification, dehumanization, deindividuation, or intolerance of other people’s views, beliefs, and identities. Speeches and articles that are shared on social media that incite division, groupism, hatred, conflict, and intolerance are classified as containing polarization. Only texts that clearly reflect attitude polarization should be classified as such. If a text includes one or more of the specified characteristics, it is classified as polarized. Conversely, social media texts that do not display any of these characteristics are classified as non-polarized. Always consider the context and the overall meaning of the text, not just individual words or phrases. INSTRUCTIONS: - Analyze the text carefully. - Classify the polarization of the given text. Use the given definition of polarization. - Provide a short reason explaining your decision. The reason should be English. - Assign a binary label: If the text is polarized, polarization is set to 1. If the text is not polarized, polarization is set to 0. - Do not include any text or formatting outside the JSON. OUTPUT FORMAT:
{"reason": "the reason of the classification decision", "polarization": 0 or 1}
Examples (if applicable):
{% if examples %}
{{examples}}
{% endif %}
Input Text:
{{user_prompt}}
F.2 Prompt for polarization type
TASK: Determine whether the following {{lang_full}} text is polarized or not polarized and which type of polarization. DEFINITION: Polarization refers to the process or phenomenon in which opinions, beliefs, or behaviors become more extreme or divided, leading to a greater distance or conflict between differing groups. Polarization denotes stereotyping, vilification, dehumanization, deindividuation, or intolerance of other people’s views, beliefs, and identities. Only texts that clearly reflect attitude polarization should be classified as such. If a text includes one or more of the specified characteristics, it is classified as polarized. Conversely, social media texts that do not display any of these characteristics are classified as non-polarized. Always consider the context and the overall meaning of the text, not just individual words or phrases. TYPES OF POLARIZATION: - Political/ideological polarization: Political polarization refers to political beliefs and affiliations becoming more extreme. - Racial or ethnic polarization: This type of polarization focuses on ethnic identity or racial origin and incites division, intolerance, and conflict between ethnic groups or races. - Religious polarization: This type of polarization focuses on religious identity and incites division, intolerance, and conflict between religious followers. - Gender/Sexual polarization: This type of polarization refers to the exclusion, discrimination, and marginalization of individuals based on their gender and sexual orientations within society, often leading to heightened tensions, misunderstandings, or conflicts. - Other: polarization texts targeting other groups/identities such as economy, technology, media, polarization, etc. INSTRUCTIONS: - Analyze the text carefully, and classify the polarization of the given text and if the text is polarized the types of polarization. - Use the given definition of polarization and their types. Provide a short reason explaining your decision. The reason should be English. - Assign a binary label for polarization: If the text is polarized, polarization is set to 1. If the text is not polarized, polarization is set to 0. - Classify the text for the different types of polarization. - The answers need to be in the following order: [political/ideological, racial/ethnic, religious, gender/sexual, other]. - The text can contain more than one type of polarization. If the text is not polarized set all polarization types to 0. Do not include any text or formatting outside the JSON. OUTPUT FORMAT:
{   "reason": "the reason of the classification decision",
    "polarization": 0 or 1,
    "polarization Types": [0/1, 0/1, 0/1, 0/1, 0/1]
{% if examples %}
**EXAMPLES**:
{{examples}}
{% endif %}
**TEXT**:
{{user_prompt}}
F.3 Prompt for polarization manifestation
TASK: Determine the manifestation type(s) of polarization expressed in the following {{lang_full}} text. LABEL SET: ["stereotype", "vilification", "dehumanization", "extreme_language", "lack_of_empathy", "invalidation"] DEFINITIONS: - stereotype: Generalizes traits to all members of a group, ignoring individual differences. - vilification: Defames or demonizes a group or person through exaggerated or biased framing. - dehumanization: Denies humanity by comparing people to animals, machines, or objects. - extreme_language: Uses absolute or hyperbolic language (e.g., “always”, “never”, “us vs. them”). - lack_of_empathy: Shows indifference or refusal to understand others’ perspectives. - invalidation: Denies or rejects the identity, legitimacy, or existence of a group or individual. INSTRUCTIONS: - Select all labels from the list that apply (one or more). - If the text is not polarized, set labels = [0]. - Provide a short reasoning sentence explaining your decision. - Do not include any text or formatting outside the JSON. OUTPUT FORMAT:
    "reason": "the reason of the classification decision",
    "polarization": 0 or 1,
    "polarization Types": [0/1, 0/1, 0/1, 0/1, 0/1, 0/1]
{% if examples %}
**EXAMPLES**:
{{examples}}
{% end if %}
**TEXT**:
{{user_prompt}}
Appendix G Error examples
Text Gold Type Reason given by model
The Cybersecurity Infrastructure Security Agency has frozen its election security work is reviewing everything it has done to help state local officials secure elections for the past 8 years. …The move accommodates Ts false claim of election fraud. @URL 0 FP The text criticizes a political figure’s claim as false and portrays a government agency as accommodating that claim, expressing a negative attitude toward the out-group and contributing to partisan division."
Four key facts Hezbollah is not an Islamic sect or denomination. It is a terrorist org founded in 1985. Hezbollah is an Iranian Shia Islamist proxy based in Lebanon. Not a Lebanese resistance movement. 0 FN The text vilifies Hezbollah by labeling it a terrorist organization and dismissing its claimed legitimacy, showing hostile stereotyping toward the group, which meets the criteria for attitude polarization.
this new optic the President use that term lightly posted on X of him as a Kingis he serious? Are we no longer a Democracy? We have an election wearly voting starting tomorrow. Im running. Did King Donny stop elections? I appreciate your courage. 1 FN The text merely questions the President’s rhetoric and expresses concern about democratic processes without demeaning, vilifying, or stereotyping any group, so it does not show attitude polarization.
Black Officers Awarded \$16M After White Colleague’s Racial Comments 1 FN The headline reports an incident involving race but does not contain demeaning language, vilification, or calls for hostility toward any group, so it is not polarized.
And thats not ethnic cleansing how ? 1 FN The brief question does not contain explicit negative or vilifying language toward any group, nor does it promote division or intolerance, so it lacks clear attitude polarization.
Table 8: Example texts, the human annotated gold labels, the error type (FN=False Negative, FP=False Positive) and the Reason given by the model
Appendix H Dataset Samples
The dataset samples are available at the following link: https://anonymous.4open.science/r/trial-data-05C6/README.md.
@@ A Federal Investigation needs to be launched against TWITTER Fraud and Election Interference P EL, I
Text Lang Polarized
@@
Text Lang Polarized